Nielsen Company reported that children between the ages of 2 and 5 watch an average of 25 hours of television per week. Assume the variable is normally distributed and the standard deviation is 3 hours. If a group of 20 children between the ages of 2 and 5 are randomly selected, find the probability that the mean of the number of hours they watch television will be greater than 26.3 hours.

 

 

 


1
Expert's answer
2021-11-22T16:42:46-0500

"z=\frac{x-\mu}{\sigma/\sqrt n}=\frac{26.3-25}{3/\sqrt{20}}=1.94"


"P(x>26.3)=1-P(z<1.94)=1-0.9738=0.262"